# Evaluating The Explainability of State-of-the-Art Deep Learning-based Network Intrusion Detection Systems

### v3

@@ Abstract
Network Intrusion Detection Systems (NIDSs) which use deep learning (DL) models achieve high detection performance and accuracy while avoiding dependence on fixed signatures extracted from attack artifacts. However, there is a noticeable hesitance among network security experts and practitioners when it comes to deploying DL-based NIDSs in real-world production environments due to their black-box nature, i.e., how and why the underlying models make their decisions. In this work, we analyze state-of-the-art DL-based NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE, SHAP) through extensive experiments with two different attack datasets. Using the explanations generated for the models’ decisions, the most prominent features used by each NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model as well as the explanations generated across the NIDS models for a given xAI method. Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). The results show that: (1) some DL-based NIDS models can be better explained than other models, (2) xAI explanations are in conflict for most of the NIDS models considered in this work and (3) some NIDS models are more vulnerable to inductive bias than other models.
Index Terms:
@@
I Introduction
@@
In this work, we analyze state-of-the-art DL-based NIDS models using explainable AI (xAI) techniques (e.g., TRUSTEE [7], SHAP [8]). Except Kitsune, other NIDS source codes do not provide trained ML models or training/test data in a format which can be directly input to the xAI methods’ APIs. Therefore, we make changes in the NIDS source codes to satisfy the APIs’ requirements. Through extensive experiments with two different attack datasets and using both global explanations and local explanations for the models’ decisions, the most prominent features used by each DL-based NIDS model considered are presented. We compare the explanations generated across xAI methods for a given NIDS model (using our proposed metric) as well as the explanations generated across the NIDS models for a given xAI method (using standard evaluation metrics). Finally, we evaluate the vulnerability of each NIDS model to inductive bias (artifacts learnt from training data). To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art DL-based NIDSs other than Kitsune using xAI techniques.
@@
  • We analyze state-of-the-art DL-based NIDS models using explainable AI techniques.
  • We present the most prominent features used by each NIDS model to make decisions on input data samples using the explanations generated earlier.
  • Using the same explanations and some evaluation metrics, we illustrate that:
    • Some DL-based NIDS models can be better explained (in terms of interpretable decision trees) than other models.
    • xAI explanations generated using different tools (e.g., TRUSTEE, SHAP) are in conflict with each other for most of the NIDS models considered.
    • Some NIDS models are more vulnerable to inductive bias (in terms of learnt packet rates) than other models.
The rest of the paper is organized as follows. In Section II, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on DL-based security systems and explainable DL-based attack detection. In Section III, we present a background on the types of explanation methods for ML models, prominent black-box xAI methods, and an overview of the state-of-the-art DL-based NIDSs to which we apply explanation techniques in this work. In Section IV, we explain the issues with applying xAI methods to the DL-based NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of xAI methods which covers the relation between explanations generated by different xAI methods and the detection of inductive bias in the DL-based NIDSs under consideration in this work. Finally, in Section V, we evaluate the explanations generated for each of the NIDSs using xAI methods, compare the explanations across the NIDS models as well across xAI methods, and discuss insights gleaned from the main results obtained thus far.
II Related Work
As a result, we can see the beginning of efforts in the research community to apply xAI techniques to ML-based security systems. A few works [11, 12, 13, 14] have employed xAI methods such as LIME [15] and SHAP [8] to assess the quality of local/global explanations generated for conventional machine learning classifiers (AdaBoost, k-nearest neighbour, Multi-layer Perceptron, Random Forest, XGBoost, Light Gradient-Boosting Machine, Gradient Boosting Classifier) applied to network intrusion datasets (NSL-KDD dataset [16], CICIDS-2017 [17]) in terms of criteria such as faithfulness, stability, complexity, and sensitivity. In [18], the authors have delved into six explanation methods, evaluating their effectiveness across four security systems described in existing literature. These systems leverage DL techniques to detect Android malware, malicious PDF files and security vulnerabilities. The assessment criteria encompass both general properties of deep learning and domain-specific aspects relevant to security.
  • Existing works that evaluate xAI methods on ML/DL-based NIDSs ([11, 12, 13, 14, 20, 22]): (1) do not compare the explainability of those NIDSs with that of state-of-the-art NIDSs and (2) do not include the latest xAI tools for ML/DL-based network security such as TRUSTEE [7].
III Background
@@
III-A Explanation Methods- Types
@@
III-B Prominent Black-box XAI Tools
@@
DL-based NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services vendor. Therefore, we assume that the DL-based NIDSs being explained in this work are available to us as a black-box system only and we use xAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box xAI tool that we have tested with as part of this work.
@@
Overview of TRUSTEE: TRUSTEE [7] is a state-of-the-art interpretability tool for ML/DL-based security solutions which generates interpretable decision trees (DT) for a given ML/DL model. It can be used to find out inductive biases in black-box models (e.g., shortcut learning, spurious correlations, sensitiveness to out-of-distribution samples). While generating DT, TRUSTEE focuses on the following desirable properties: model-agnosticism (should be applicable to any given black-box model), model fidelity (how accurate the DT is with respect to the original ML model), model stability (the decision rules should correctly describe how the given black-box model makes a significant number of its decisions and should be largely insensitive to the particular data samples that TRUSTEE used in the process of selecting its final DT explanation) and model complexity (decision rules should be readily recognizable by domain experts). The important aspects of generated DT explanation are summarized in a trust report which helps end users determine whether the given black-box model suffers from the problem of underspecification and can be trusted.
@@
Overview of LIME: Ribeiro et al. [15] introduced LIME, one of the first black-box methods for explaining neural networks that is further extended by SHAP [8]. Both works are motivated by the need to allow users to be able to trust ML models as well as their predictions. LIME is designed to explain the predictions of any classifier or regressor in a faithful way, by approximating it locally with an interpretable model. The main characteristics of LIME are:
  • Interpretable, i.e., it provides qualitative understanding of the relation between the input variables and the response.
  • Locally faithful, i.e., it corresponds to how the model behaves in the vicinity of the instance being predicted.
  • Model-agnostic, i.e., it treats the original ML model as a black box.
LIME aims at approximating the decision function, $f$ of the model being explained by solving the following optimization problem:
@@
xAI Evaluation Metrics: Several metrics have been proposed in literature to quantify the effectiveness of xAI techniques. The authors in [27] have categorized existing XAI evaluation metrics into subjective and objective metrics. Subjective metrics rely on human feedback, e.g., comprehensibility, satisfaction, and persuasiveness. This feedback can come from randomly selected users or domain experts. Objective metrics are grounded in formal definitions, e.g., explanation completeness which quantifies the extent to which a prediction’s contributing features are captured, and perturbation metrics which assess sensitivity to input changes by comparing outputs and considering model parameter variations across different instances of the same model.
III-C State-of-the-art Deep Learning-based NIDSs
@@
The above state-of-the-art systems were selected since: (1) they were published in prominent avenues (NDSS, USENIX Security, ACSAC) and include works published as recently as 2023, (2) provide a clear performance comparison with Kitsune (which is the most cited deep learning-based NIDS with an open-source implementation) and (3) they cover a diverse ground on the application of deep learning in network intrusion detection. Kitsune uses an ensemble of autoencoders while ENIDrift builds on top of that by including a sub-classifier generation and updation module to address concept drift in real-world networks. HorusEye’s Magnifier switches the ensemble of autoencoders for a single asymmetric autoencoder to keep the solution lightweight and reduce the false-positive rate. HELAD uses a single autoencoder as well but combines it with an LSTM network.
IV xAI Analysis of Deep Learning-based NIDSs
In this work, we analyze four state-of-the-art DL-based NIDSs (Kitsune, HorusEeye, ENIDrift, HELAD) using xAI techniques. However, xAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the NIDSs under consideration. This is because xAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with xAI methods. Still, there were a few xAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code. Additionally, we did not limit ourselves to a direct application of xAI methods to the NIDSs but we also investigated: (1) the relation between the explanations generated for a given NIDS by different xAI methods, and (2) the presence of inductive biases in the NIDS implementations.
IV-A xAI-compatible NIDS Implementation
@@
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML/DL models which do not have $predict\_proba()$ function implemented. Since all the state-of-the-art DL-based NIDSs considered in this paper do not implement $predict\_proba()$ function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, $predict\_proba()$ function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML/DL models such as the NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP.
IV-B Investigation Beyond Direct xAI Analysis
Relation between TRUSTEE and SHAP Interpretations: To establish the relationship between TRUSTEE global explanations and the SHAP local explanations of the black-box model underlying a given NIDS, we can follow two approaches:
@@ IV-B Investigation Beyond Direct xAI Analysis
Under the systematic approach, for a given subset of data samples $D$, if $T$ is the set of features used in the selected root node-leaf node DT path decisions and $S$ is the set of features identified by kernel SHAP that contribute most to the classification decision for majority of the samples in $D$, then $T\subset S$ should hold true for the TRUSTEE DT explanation to agree with the SHAP explanation of the black-box model at the subset level. To form the set of features $S$, the mean absolute value of kernel SHAP values for multiple data samples are calculated for every feature, and subsequently the features are ranked according to their aggregate SHAP values to find the top features. For a black-box NIDS model, the local explanation (SHAP) may agree with the global explanation (TRUSTEE) depending on the data subset on which they are executed. However, not all local explanations may agree with the global explanation. To evaluate the level of agreement between TRUSTEE and SHAP for an NIDS trained on a given dataset, we propose the following approach:
@@
We also define the m-point average TRUSTEE-SHAP agreement score for an NIDS as the average of TRUSTEE-SHAP agreement scores for $m$ different data subsets. It can be expressed as:
@@
Inductive Bias: A number of ML/DL models face the issue of underspecification. This term refers to the challenge of determining whether a model’s high performance is due to its genuine ability to capture essential patterns in the data or merely the result of certain inductive biases it has learned. These biases often appear as spurious correlations, a lack of ability to adapt to data samples with variable statistics, shortcut learning strategies, etc. The presence of such biases in trained models undermines their reliability and expected performance in real-world applications. Therefore, identifying these inductive biases is crucial for establishing trust in ML/DL models. To detect the presence of an inductive bias in the DL-based NIDSs under consideration in this work, such as vulnerability to variable packet rate samples, we use tampered malicious traffic traces [7] which alter the attack segments of the original traffic traces by distributing the malicious packets in such a way that the packet rate remains within a specified limit.
V Experimental Results
In this section, we present explanations of state-of-the-art DL-based NIDSs generated using xAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and xAI method, we wrote Python scripts for generating the explanation using the xAI method’s API and the modified source code of the NIDS (to make it compatible with the xAI method’s API implementation). Using the generated explanations, we present the most prominent features used by each NIDS model to make decisions on input data samples. We also investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP by using the systematic approach for comparing TRUSTEE and SHAP explanations at the DT subset-level as explained in sub-section IV-B. Subsequently, we investigate the presence of inductive biases (e.g., vulnerability to variable packet rate samples) in the NIDS implementations using tampered malicious traffic traces as explained in sub-section IV-B. Finally, we compare the explanations generated across the four NIDSs by TRUSTEE using the fidelity metric and by kernel SHAP using general xAI evaluation metrics.

### v4

# Evaluating The Explainability of Deep Learning-based Network Intrusion Detection Systems

@@ Abstract
State-of-the-art deep learning (DL)-based network intrusion detection systems (NIDSs) offer limited “explainability”. For example, how do they make their decisions? Do they suffer from hidden correlations? Prior works have applied eXplainable AI (XAI) techniques to ML-based security systems such as conventional ML classifiers trained on public network intrusion datasets, Android malware detection and malicious PDF file detection. However, those works have not evaluated XAI methods on state-of-the-art DL-based NIDSs and do not use latest XAI tools. In this work, we analyze state-of-the-art DL-based NIDS models using conventional as well as recently proposed XAI techniques through extensive experiments with different attack datasets. Furthermore, we introduce a criteria to evaluate the level of agreement between global- and local-level explanations generated for an NIDS. Using this criteria in addition to other security-focused criteria, we compare the explanations generated across XAI methods. The results show that: (1) the decisions of some DL-based NIDS models can be better explained than other models, (2) XAI explanations generated using different tools are in conflict for most of the NIDS models considered in this work and (3) there are significant differences between XAI methods in terms of some security-focused criteria. Based on our results, we make recommendations on how to achieve a balance between explainability and model detection performance.
keywords:
@@
[…]
organization=Cyber Security Strategic Technology Centre, ST Engineering,addressline=600 W Camp Rd, city=Singapore, postcode=797654, state=Singapore, country=Singapore
1 Introduction
@@
In this work, we analyze state-of-the-art DL-based NIDS models using eXplainable AI (XAI) techniques (e.g., TRUSTEE [7], SHAP [8]). In addition, we develop a criteria to evaluate the level of agreement between global- and local-level explanations generated for an NIDS. We compare the explanations generated across XAI methods using our criteria and other security-focused evaluation criteria. To the best of our knowledge, our work is the first one to generate explanations for state-of-the-art DL-based NIDSs other than Kitsune using XAI techniques. Though a few works [9, 10, 11, 12, 13] have applied XAI techniques to ML-based security systems (conventional ML classifiers trained on public network intrusion datasets, Android malware detection, malicious PDF file detection) in the past, none of those systems corresponded to state-of-the-art DL-based NIDSs.
@@
  • We analyze state-of-the-art DL-based NIDS models using XAI techniques. The application of XAI techniques is not trivial and we make several modifications to the NIDS implementations.
  • We propose a criteria to evaluate the level of agreement between global- and local-level explanations generated for an NIDS.
  • We compare the explanations generated across XAI methods using our criteria and other security-focused evaluation criteria. This highlights the need (for a security practitioner) to select the best method for a particular NIDS as per the preferred criteria.
  • We propose an approach (for a security practitioner) to select the best DL-based NIDS for a specific use-case by achieving the desired balance between detection performance and explainability.
The rest of the paper is organized as follows. In Section 2, we present a background on the types of explanation methods for ML models, prominent black-box XAI methods, and an overview of the state-of-the-art DL-based NIDSs to which we apply explanation techniques in this work. In Section 3, we explain the issues with applying XAI methods to the DL-based NIDSs directly using their open-source implementations, our approach to addressing those issues and details of our investigation beyond direct application of XAI methods which covers the relation between explanations generated by different XAI methods. In Section 4, we evaluate the explanations generated for each of the NIDSs using XAI methods, compare the explanations across XAI methods, and in Section 5, we discuss insights gleaned from the main results obtained thus far. In Section 6, we review published research on the need for explanation of ML systems targeted at cybersecurity, evaluation of explanation methods on DL-based security systems and explainable DL-based attack detection. Finally, we summarise the key results obtained and provide future research directions in Section 7.
2 Background
@@
2.1 Explanation Methods- Types
@@
2.2 Prominent Black-box XAI Tools
@@
DL-based NIDS products are expected to be used by end users/customers who do not have access to the ML model and its parameters underlying the NIDS. This is to maintain confidentiality of the NIDS implementation by the cybersecurity services provider. Therefore, we assume that the DL-based NIDSs being explained in this work are available to us as a black-box system only and we use XAI techniques designed to explain such black-box ML systems. In what follows, we present a brief overview of each black-box XAI tool that we have tested with as part of this work.
@@
Overview of TRUSTEE: TRUSTEE [7] is a state-of-the-art interpretability tool for ML/DL-based security solutions which generates interpretable decision trees (DT) for a given ML/DL model. It can be used to find out inductive biases in black-box models (e.g., shortcut learning, spurious correlations). While generating DT, TRUSTEE focuses on the following desirable properties: model-agnosticism (should be applicable to any given black-box model), model fidelity (how accurate the DT is with respect to the original ML model), model stability (the decision rules should correctly describe how the given black-box model makes a significant number of its decisions and should be largely insensitive to the particular data samples that TRUSTEE used in the process of selecting its final DT explanation) and model complexity (decision rules should be readily recognizable by domain experts). The important aspects of generated DT explanation are summarized in a trust report which helps end users determine whether the given black-box model suffers from the problem of underspecification and can be trusted.
In the TRUSTEE algorithm, there is an inner loop which is designed to generate different high fidelity DT explanations, one per iteration. It does so by applying a teacher-student dynamic derived from imitation learning that uses the blackbox model, $\pi^{*}$ as an oracle in conjunction with a carefully curated dataset $D$’ to guide the training of a surrogate “white-box” model in the form of a DT that imitates the black-box’s decisions. In contrast, the purpose of the outer loop is (i) to select from among the $N$ high fidelity DTs that have been generated in the process of executing the inner loop the DT with the highest fidelity, (ii) to transform this resulting DT into a high-fidelity and low-complexity DT by means of a post-processing step that consists of applying a purposefully developed pruning method called top-$k$ pruning which carefully scrutinizes only the top $k$ branches of an extracted high-fidelity DT, ranked by the number of input samples a branch classifies, and (iii) to consider all $S$ high-fidelity and low-complexity DTs that have been generated in the process of executing the outer loop and output the one that is the most stable in the sense of having the highest mean agreement among these $S$ DTs.
@@
Overview of LIME: Ribeiro et al. [14] introduced LIME, one of the first black-box methods for explaining neural networks that is further extended by SHAP [8]. Both works are motivated by the need to allow users to be able to trust ML models as well as their predictions. LIME is designed to explain the predictions of any classifier or regressor in a faithful way, by approximating it locally with an interpretable model. LIME aims at approximating the decision function, $f$ of the model being explained by solving the following optimization problem:
@@
In order to learn the local behavior of $f$ as the interpretable inputs vary, LIME approximates $\mathcal{L}(f,g,\pi_{x})$ by drawing samples, weighted by $\pi_{x}$. The instances around $x$ are sampled by drawing nonzero elements of $x$ uniformly at random. Next, LIME creates a series of $l$ perturbations of $x$, denoted as $\tilde{x_{1}}$, $\dots$, $\tilde{x_{l}}$ by setting entries in the vector $x$ to 0 randomly. The method then proceeds by predicting a label $f(\tilde{x_{i}})=\tilde{y_{i}}$ for each $\tilde{x_{i}}$ of the $l$ perturbations. This strategy enables the method to approximate the local neighborhood of $f$ at the point $f(x)$. LIME approximates the decision boundary by a weighted linear regression model,
$\mathcal{L}(f,g,\pi_{x})=\sum_{i=1}^{l}\pi_{x}(\tilde{x_{i}})(f(\tilde{x_{i}})-g(\tilde{x_{i}}))^{2},$ (2)
where $\mathcal{G}$ is the set of all linear functions and $\pi_{x}$ is a function indicating the difference between the input $x$ and a perturbation $\tilde{x}$.
@@
2.3 State-of-the-art Deep Learning-based NIDSs
@@
Deep ResNIDS [20] uses a combination of three deep-neural network stages to detect novel attacks at the packet traffic-level: a malicious packet detector (supervised) which detects known malicious traffic, an anomaly detector (unsupervised) which distinguishes novel packets (out-of-distribution, zero-day and adversarial attacks) from truly benign packets and a novelty detector (supervised) which distinguishes adversarial manipulated packets from truly novel packets. To re-train the models in above three stages for addressing evolving attacks, a transfer learning approach is used wherein all but few of the trained layers’ weights are frozen and an additional layer is appended to the model for training.
The above state-of-the-art systems were selected since: (1) they were published in prominent avenues (NDSS, USENIX Security, ACSAC, Computers & Security journal) and include works published as recently as 2024, (2) they provide a clear performance comparison with Kitsune (which is the most cited deep learning-based NIDS with an open-source implementation) and (3) they cover a diverse ground on the application of deep learning in network intrusion detection. Kitsune uses an ensemble of autoencoders while ENIDrift builds on top of that by including a sub-classifier generation and updation module to address concept drift in real-world networks. HorusEye’s Magnifier switches the ensemble of autoencoders for a single asymmetric autoencoder to keep the solution lightweight and reduce the false-positive rate. HELAD uses a single autoencoder as well but combines it with an LSTM network. Deep ResNIDS employs three stages of deep-learning classifiers to detect novel attacks out of which the latter two stages (anomaly detector and novelty detector) use autoencoders.
3 XAI Evaluation of Deep Learning-based NIDSs
In this work, we analyze state-of-the-art DL-based NIDSs presented in previous section using XAI techniques. However, XAI techniques such as TRUSTEE and kernel SHAP can not be applied directly to the NIDSs under consideration. This is because XAI methods have certain requirements at the API level which are not fulfilled by the current implementations of the NIDSs. We had to modify the NIDS implementations to make them compatible with XAI methods. Still, there were a few XAI methods whose requirements were found to be incompatible with the NIDS implementations unless we made major changes to the source code.
3.1 XAI-compatible NIDS Implementation
@@
While running LIME analysis of above NIDSs, we received an error that LIME does not support ML/DL models which do not have predict_proba()predict\_proba() function implemented. Since all the state-of-the-art DL-based NIDSs considered in this paper do not implement predict_proba()predict\_proba() function, we are forced to leave out LIME analysis of the NIDSs from this paper. Here, predict_proba()predict\_proba() function returns the predicted probabilities of the input features belonging to each category. This method, instead of returning a discrete class, returns the probabilities associated with each class. This is useful when we not only want to know the predicted category of the input features, but also the model’s confidence in its prediction. Since SHAP is a generalized version of LIME, excluding LIME from our analysis would not affect our results substantially.
While running LEMNA analysis of above NIDSs, we found that the current implementation (last changed 5 years ago) only supports basic MLP, CNN and RNN models. Custom ML/DL models such as the NIDS models which are under consideration in this work are not supported. Further, LEMNA does not output feature-based model explanations for processing by cybersecurity experts such as those offered by TRUSTEE and SHAP. Though LEMNA improves upon LIME by using non-linear local approximation and handling feature dependencies, it is still inferior to SHAP due to the latter’s high mathematical rigor and strong guarantees such as consistency and local accuracy. Thus, excluding LEMNA from our analysis would not affect our results substantially.
3.2 Evaluation Criteria
To evaluate the performance of various explanation methods on DL-based NIDSs, we propose a new criteria called TRUSTEE-SHAP Agreement Score. Further, we use the above criteria as well as security-related criteria introduced in [21] (Stability and Adversarial Robustness) to compare the explanation methods. These criteria should come in handy for security practitioners while selecting the best explanation method suited to their use-case.
3.2.1 TRUSTEE-SHAP Agreement Score
To establish the relationship between TRUSTEE global explanations and the SHAP local explanations of the black-box model underlying a given NIDS, we can follow two approaches:
@@ 3.2.1 TRUSTEE-SHAP Agreement Score
Under the systematic approach, for a given subset of data samples $D$, if $T$ is the set of features used in the selected root node-leaf node DT path decisions and $S$ is the set of features identified by kernel SHAP that contribute most to the classification decision for majority of the samples in $D$, then $T\subset S$ should hold true for the TRUSTEE DT explanation to agree with the SHAP explanation of the black-box model at the subset level. To form the set of features $S$, the mean absolute value of kernel SHAP values for multiple data samples are calculated for every feature, and subsequently the features are ranked according to their aggregate SHAP values to find the top features. For a black-box NIDS model, the local explanation (SHAP) may agree with the global explanation (TRUSTEE) depending on the data subset on which they are executed. However, not all local explanations may agree with the global explanation.
To evaluate the level of agreement between TRUSTEE and SHAP for an NIDS trained on a given dataset, we propose the following approach:
@@
We define the m-point average TRUSTEE-SHAP agreement score for an NIDS as the average of TRUSTEE-SHAP agreement scores for $m$ different data subsets. It can be expressed as:
@@
3.2.2 Adversarial Robustness
Multiple works [22, 23, 24] have shown that small input perturbations can lead to significant changes in explanations while the classification output remains same. Another class of adversarial attacks against LIME and SHAP [25, 26] replaces a black-box model (with biased predictions) with a surrogate model to hide the bias (e.g., from external auditors). A user querying the model service for the explanation of a decision by the black-box model is instead provided the explanations of the surrogate model. Though defensive strategies to increase the robustness of explanation techniques have been proposed in recent works, there is still a lack of formal analysis of that robustness [27]. It should be noted that all of the above works are focused on image classification. Adversarial perturbations for NIDS models may not correspond to valid traffic, and even if they do, they may not lead to significant changes in explanation. We evaluate the robustness of the explanation generated for a black-box NIDS model using informal techniques inspired by literature.
3.2.3 Stability
The explanations generated for a black-box model using a particular explanation method over multiple runs may not match completely, i.e., some features which are part an explanation generated for one run may not appear in the explanation generated in another run. For a security practitioner, if the feature-based explanation is not stable over multiple runs, they may decide against using it. As per [21], to determine the stability of an explanation method, we calculate the Intersection Size (IS) for runs $i$ and $j$ as:
$IS(i,j)=\frac{|T_{i}\cap T_{j}|}{k},$ (6)
where $T_{i}$ and $T_{j}$ are the sets of top $k$ relevant features from the explanation generated in runs $i$ and $j$, respectively. An explanation method is defined as stable if the corresponding intersection size is close to $1$, i.e., $IS(i,j)>1-\epsilon$ for a small threshold $\epsilon$.
4 Experimental Results
In this section, we present explanations of state-of-the-art DL-based NIDSs generated using XAI methods such as TRUSTEE and kernel SHAP. For each pair of NIDS and XAI method, we wrote Python scripts for generating the explanation using the XAI method’s API and the modified source code of the NIDS (to make it compatible with the XAI method’s API implementation). Using the generated explanations and further analysis, we aim to answer the following research questions:
  1. 1.
    RQ1: How do TRUSTEE and SHAP compare against each other in terms of various evaluation criteria?
    We use the TRUSTEE-SHAP agreement score introduced in sub-section 3.2 to investigate the relation between the explanations generated for a given NIDS by TRUSTEE and kernel SHAP. We also use security-focused criteria such as adversarial robustness and stability to compare TRUSTEE versus SHAP explanations.
  2. 2.
    RQ2: What are the most prominent features used by each NIDS model to make decisions on input data samples?
    The explanations generated for each NIDS model by TRUSTEE and kernel SHAP rank the features used for making decisions on input data samples. TRUSTEE ranks the features based on the percentage of total number of test data samples classified using those features while kernel SHAP ranks the features based on the mean absolute SHAP values for those features.